# MsgABD: EveryShipout initializing macros

### v1

@@ 3.2 Nonlinear MPNNs
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.11)
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\hat{\bm{A}}^{\ell}\right)_{ij}$ (3.12)
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ $\leq\zeta_{2}^{\left(L\right)}\left(\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[I_{N}+\sum_{\ell=1}^{L}\prod_{k=1}^{\ell}\hat{\bm{A}}^{\left(k\right)}\right]\right)_{ij}$ (3.13)
$=\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$ (3.14)
Figure 5 shows the plot of the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$ (i.e. for DropEdge), as in the upper bound above, with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$. We observe that the sensitivity between nearby nodes marginally increases, while that between distant nodes notably decreases (similar to Figure 1), suggesting significant over-squashing. Similar observations can be made with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, and for other DropEdge-variants.
Source-and-Target Message Functions. Topping et al. (2022, Lemma 1) showed that if the aggregation function is instead given by
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.15)
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{3}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.16)
With random edge-dropping, this bound can be adapted as follows:
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.17)
which is similar to Equation 3.10, only with a different proportionality constant, that is anyway independent of the choice of nodes. Here, again, we invoke Theorem 3.1 to conclude that $(\mathbb{E}[\hat{\bm{A}}]^{L})_{ij}$ decreases monotonically with increasing DropEdge probability $q$. This implies that, in a non-linear MPNN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, DropEdge lowers the sensitivity bound given above. Empirical results in Figure 4 support the same conclusion for $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$.
Message of the Section: Studying the expected propagation matrix allows us to predict the effect of random edge-dropping methods on information propagation in MPNNs. Specifically, DropEdge, DropAgg and DropGNN increase the sensitivity of nodes to their neighbors, but decrease it to nodes farther off. This suggests that these methods would be unsuitable for long-range tasks, where it is imperative to facilitate communication between distant nodes.
@@ 4 Experiments
Our theoretical analysis indicates that DropEdge may degrade the performance of GNNs on tasks that depend on capturing LRIs. In this section, we test this hypothesis by evaluating DropEdge models on both short-range and long-range tasks.
4.1 Setup
@@
Datasets. Although identifying whether a task requires modeling LRIs can be challenging, understanding the structure of the datasets can provide some insight. For example, homophilic datasets have local consistency in node labels, i.e. nodes closely connected to each other have similar labels. On the other hand, in heterophilic datasets, nearby nodes often have dissimilar labels. Since DropEdge-variants increase the sensitivity of a node’s representations to its immediate neighbors, and reduce its sensitivity to distant nodes, we expect it to improve performance on homophilic datasets but harm performance on heterophilic ones (see Topping et al. (2022)). In this work, we use Cora (McCallum et al., 2000) and CiteSeer (Giles et al., 1998) as representatives of homophilic datasets (Zhu et al., 2020), and Squirrel, Chameleon and TwitchDE (Rozemberczki et al., 2021) to represent heterophilic datasets (Lim et al., 2021). The networks’ statistics are presented in Table 1, where we can note the significantly lower homophily measures of heterophilic datasets.
Dataset Nodes Edges Features Classes Homophily
Cora 2,708 10,556 1,433 7 0.766
CiteSeer 3,327 9,104 3,703 6 0.627
Chameleon 2,277 36,051 2,325 5 0.062
Squirrel 5,201 216,933 2,089 5 0.025
TwitchDE 9,498 306,276 128 2 0.142
Table 1: Dataset statistics. Number of edges excludes self-loops. Homophily measures from Lim et al. (2021).
Models. We use two MPNN architectures, GCN and GAT, to demonstrate the effect of DropEdge. GCN satisfies the model assumptions made in all the theoretical results presented in Section 3, while GAT does not fit into any of them. Therefore, GCN and GAT together provide a broad representation of different MPNN architectures. For GAT, we use 2 attentions heads in order to keep the computational load manageable, while at the same time harnessing the expressiveness of the multi-headed self-attention mechanism. As for the model depth, we vary it as $L=2,4,6,8$.
Dropout Methods. In addition to the DropEdge and its variants – DropNode, DropAgg and DropGNN – discussed in Section 3, we evaluate related stochastic regularization methods like regular dropout (Srivastava et al., 2014) (referred as Dropout), DropMessage (Fang et al., 2023) and SkipNode Lu et al. (2024). This allows us to study the pervasiveness of the problem of over-squashing over a large set of methods designed for alleviating over-smoothing.
Dropping Probability. For all the methods, the dropping probabilities are varied as $q=0.0,0.1,\ldots,0.9$, so as to reliably capture the trends in model accuracy. We adopt the common practice of turning the methods off at test-time ($q=0$), isolating their effect on optimization and generalization, which our theory does not address.
Other experimental details are reported in Appendix E. We conduct 5 independent runs for each dataset--model--dropout--probability configuration and report the average test accuracy (with early-stopping using the validation set).
4.2 Results
Dropout GNN Homophilic Heterophilic
Cora CiteSeer Chameleon Squirrel TwitchDE
DropEdge GCN $+0.396$ $+0.448$ $-0.723$ $-0.614$ $-0.509$
GAT $+0.547$ $+0.548$ $-0.236$ $-0.641$ $-0.390$
DropNode GCN $-0.320$ $-0.662$ $-0.863$ $-0.790$ $-0.465$
GAT $-0.706$ $-0.735$ $-0.855$ $-0.286$ $-0.512$
DropAgg GCN $-0.015$ $+0.294$ $-0.334$ $-0.317$ $-0.566$
GAT $-0.019$ $+0.392$ $-0.353$ $-0.398$ $-0.022$
DropGNN GCN $+0.360$ $+0.468$ $-0.746$ $-0.600$ $-0.671$
GAT $+0.236$ $+0.525$ $-0.507$ $-0.304$ $-0.139$
Dropout GCN $+0.034$ $-0.309$ $-0.610$ $-0.465$ $-0.300$
GAT $+0.120$ $-0.110$ $-0.836$ $-0.545$ $-0.397$
DropMessage GCN $-0.220$ $-0.607$ $-0.404$ $-0.364$ $-0.199$
GAT $+0.028$ $-0.368$ $-0.530$ $-0.697$ $-0.148$
SkipNode GCN $+0.161$ $+0.017$ $-0.528$ $-0.285$ $-0.244$
GAT $+0.403$ $+0.466$ $-0.436$ $+0.248$ $+0.278$
Table 2: Spearman correlation between dropping probability and test accuracy. Note the positive correlations for homophilic datasets and negative correlations for heterophilic datasets. Results suggest that random edge-dropping, although effective at improving generalization in short-range tasks, is unsuitable for long-range ones.
In Table 2, we report the rank correlation between the dropping probability and test accuracy of different dataset--model--dropout combinations. The statistics are computed over $5010\ (q=0.0,0.1,\ldots,0.9)\times 5\ (\text{samples})=$ 50 runs for each $L=2,4,6,8$, and then averaged. It is clear to see that in most combinations with the homophilic datasets Cora and CiteSeer, the correlation is positive, indicating that DropEdge and its variants improve test-time performance at short-range tasks. On the other hand, with the heterophilic datasets Chameleon, Squirrel and TwitchDE, the correlation values are negative. This suggests that the edge-dropping methods harm generalization in long-range tasks by forcing the model to overfit to short-range signals.
(a) Homophilic datasets
(b) Heterophilic datasets
Figure 2 allows us to visualize these trends for the homophilic and heterophilic datasets. For conciseness, we only display the results for DropEdge-GCNs. Clearly, on Cora and CiteSeer, the model’s test-time performance improves with increasing dropping probability, as has been reported earlier. However, on Chameleon, Squirrel and TwitchDE, the performance degrades with increasing dropping probability, as was suggested by our theoretical results. This highlights an important gap in our understanding of dropout methods – while their positive effects on model performance have been well-studied, making them popular choices for training deep GNNs, their evaluation has been limited to short-range tasks, leaving their negative impact on capturing LRIs overlooked.
Remark on DropNode. In Equation 3.4, we noted that DropNode does not suffer from loss in sensitivity. However, note that those results were in expectation. Moreover, our analysis did not account for the effects on the learning trajectory. In practice, a high DropNode probability would make it hard for information in the node features to reach distant nodes. This would prevent the model from learning to effectively combine information from large neighborhoods, harming generalization. In fact, in Table 2, we can see that it is the only dropping method with a negative correlation between dropping probability and test accuracy for each dataset--model combination, including homophilic ones. See Figure 6 for a visualization of its negative effects on test accuracy.
4.3 Over-squashing or Under-fitting?
Figure 3: DropEdge probability versus training accuracy of GCNs. The training performance improves with $q$, suggesting that the models are not underfitting. Instead, the reason for poor test-time performance (2(a)) is that models are over-fitting to short-range signals during training, resulting in poor generalization.
The results in the previous subsection suggest that using random edge-dropping to regularize model training leads to poor test-time performance. We hypothesize that this occurs because the models struggle to propagate information over long distances, causing overfitting of node representations to local neighborhoods. However, a confounding factor is at play: DropEdge variants reduce the generalization gap by preventing overfitting to the training set, which manifests as reduced training performance. If this regularization is too strong, it could lead to underfitting, which could also explain the poor test-time performance on heterophilic datasets. This concern is particularly relevant because the heterophilic networks are much larger than homophilic ones (Table 1), making them more prone to underfitting. To investigate this, we plot the training accuracies of deep DropEdge-GCNs on the heterophilic datasets; Figure 3 shows the results. It is clear that the models do not underfit as the dropping probability increases. In fact, somewhat unexpectedly, the training metrics improve. Together with the results in Figure 2, we conclude that DropEdge-like methods are detrimental in long-range tasks, as they cause overfitting to short-range artifacts in the training data, resulting in poor generalization at test-time.
Message of the Section: Random edge-dropping algorithms reduce the receptive field of MPNNs, forcing them to fit to short-range signals in the training set. While this may make deep GNNs suitable for homophilic datasets, it results in overfitting on heterophilic datasets, which leads to poor test-time performance. Therefore, these methods should be used only after carefully understanding the task at hand.
@@ 5 Conclusion
Our analysis points out a key assumption in algorithms designed for training deep GNNs: the idea that if a deep GNN is trainable, it must be able to model LRIs. Our results suggest that this, in fact, need not be true – we theoretically and empirically show that DropEdge-like algorithms exacerbate the over-squashing problem in deep GNNs, and degrade their performance on long-range tasks. Our results highlight a need for a thorough evaluation of methods employed when training deep GNNs, with regards to their capacity to capture LRIs. This will allow us to reliably deploy them in real-life, since we can be assured that the models did not simply overfit to short-range signals.
Limitations. While our theoretical analysis successfully predicts how DropEdge-variants affect test performance on short-range and long-range tasks, it is based on several simplifying assumptions. These assumptions, although standard in the literature, limit the generalizability of our conclusions to other architectures. Specifically, our analysis focuses on certain classes of MPNNs, excluding several GNN architectures specifically designed to enhance long-distance information propagation (see Singh (2024) for a review).
Practical Considerations. Previous studies have shown that random edge-dropping algorithms can effectively enhance generalization performance in short-range tasks, and our findings support this conclusion. However, we have also demonstrated that such algorithms can negatively impact over-squashing in MPNNs and harm test-time performance in long-range tasks. Therefore, we recommend exercising caution when using such methods, as careless application can result in models that generalize poorly, which can be detrimental in critical applications.
Future Directions. This work focuses on node-classification tasks, but it is also important to understand the effect of random edge-dropping in other practical settings, e.g. link prediction, and even graph-level tasks. In general, there is a need for a broader investigation into methods designed for training deep GNNs. Specifically, analyzing various strategies designed for mitigating over-smoothing (see Rusch et al. (2023) for a review), particularly in the context of over-squashing, could be invaluable for designing deep GNNs for long-range tasks.

### v2

@@ 4 Sensitivity-Aware DropEdge
$\begin{split}c&=\frac{\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}_{ij}]}{\mathbb{E}_{\mathsf{ND}}[\hat{\bm{A}}_{ij}]}=\frac{d_{i}+1}{d_{i}}\left(1-\frac{1-q_{i}^{d_{i}+1}}{\left(1-q_{i}\right)\left(d_{i}+1\right)}\right)\Longrightarrow 1-c=\frac{q_{i}-q_{i}^{d_{i}+1}}{d_{i}\left(1-q_{i}\right)}\end{split}$ (4.1)
where $c$ is the fraction of information preserved, e.g. 95%. We can solve for $q_{i}$ and mask the incoming edges to node $i$ accordingly; we name this algorithm DropSens. In Section E.3, we present a Python implementation of the algorithm, as well as a computationally efficient approximation to Equation 4.1. In 1(b), we observe that DropSens improves sensitivity between distant nodes, compared to DropEdge.
@@ 5 Experiments
Our theoretical analysis indicates that random dropping may degrade the performance of GNNs in tasks that depend on capturing LRIs. In this section, we test this hypothesis by evaluating DropEdge- and Dropout-variants on both synthetic and real-world datasets. A complete description of the datasets is provided in Section E.1, and the experimental details are in Section E.2.
5.1 Synthetic Datasets
Figure 2: Train and test MAE of 11-layer GCNs on the SyntheticZINC dataset, averaged over 10 initializations.
SyntheticZINC [33] is a synthetic variant of the ZINC dataset [42], designed to study the effect of information mixing in graph learning. Node features are sparsely assigned, and the target requires non-linear mixing of two selected nodes’ features, chosen based on their commute time [10]. We vary the mixing level and evaluate an 11-layer GCN, ensuring sufficient message passing. For better readability, we only test DropEdge, Dropout and DropMessage – the three more popular methods used for training deep GNNs.
The results are presented in Figure 2, where we can observe that the mean absolute error (MAE) increases with the commute time percentile used to select the node pairs, as was hypothesized and evidenced in Giovanni et al. [33]. Additionally, we observe that both train and test performance decline when using dropout with a probability as low as 0.2, and even more so with a higher probability of 0.5. These results provide strong evidence for the detrimental effects of dropout methods in modelling long-range interactions, supporting our theoretical analysis.
5.2 Real-world Datasets
Table 1: Difference in mean test accuracy (%) between the best performing configuration of each dropout method and the baseline NoDrop model. Cell colors represent p-values from a t-test evaluating whether dropout improves performance: green indicates significance at 90% confidence, while red denotes insignificant results.
(a) Node classification tasks.
GNN Dropout Homophilic Networks Heterophilic Networks
Cora CiteSeer PubMed Chameleon Squirrel TwitchDE
GCN DropEdge $+0.419$ $+0.686$ $+0.385$ $-0.634$ $+0.009$ $-0.093$
DropNode $+0.373$ $+0.197$ $+0.841$ $-0.674$ $-0.656$ $-0.113$
DropAgg $+0.224$ $+0.486$ $-0.245$ $-10.640$ $-13.970$ $-6.674$
DropGNN $+0.396$ $+0.820$ $+0.506$ $-1.774$ $-0.291$ $-0.395$
Dropout $+0.628$ $+0.128$ $+1.218$ $-1.395$ $-0.110$ $-0.260$
DropMessage $+0.030$ $-0.388$ $+1.217$ $+1.322$ $+0.320$ $+0.160$
GIN DropEdge $-0.276$ $-0.529$ $-0.034$ $-0.726$ $+0.289$ $-0.124$
DropNode $+0.379$ $+0.926$ $+0.296$ $-1.832$ $-0.265$ $-0.171$
DropAgg $+0.095$ $-0.254$ $-0.487$ $-1.153$ $+0.275$ $+0.205$
DropGNN $-0.478$ $-2.284$ $-1.366$ $-1.642$ $-0.066$ $-0.008$
Dropout $+1.347$ $+0.011$ $+0.368$ $-2.702$ $-0.152$ $-0.573$
DropMessage $+2.602$ $+0.219$ $+1.030$ $+0.943$ $+0.108$ $-0.025$
(b) Graph classification tasks.
GNN Dropout Molecular Networks Social Networks
Mutag Proteins Enzymes Reddit IMDb Collab
GCN DropEdge $-1.100$ $+1.750$ $-0.589$ $-8.380$ $+1.300$ $-1.145$
DropNode $-5.100$ $+2.339$ $-2.292$ $-7.440$ $+2.720$ $-4.054$
DropAgg $-0.900$ $+1.214$ $-0.553$ $-16.460$ $+3.580$ $-30.386$
DropGNN $-0.200$ $+1.589$ $-2.918$ $-11.860$ $+1.540$ $+0.705$
Dropout $-0.300$ $+2.018$ $-6.208$ $-6.050$ $+1.240$ $-1.729$
DropMessage $+2.000$ $+2.143$ $-4.906$ $-7.360$ $+1.300$ $-0.330$
GIN DropEdge $-1.100$ $-1.804$ $-4.716$ $-2.700$ $-1.820$ $-0.642$
DropNode $-3.800$ $-2.911$ $-0.493$ $+0.550$ $-0.120$ $+1.206$
DropAgg $-3.800$ $-1.982$ $-0.640$ $+0.930$ $-1.640$ $-5.943$
DropGNN $-7.000$ $-2.750$ $-3.694$ $+1.210$ $-5.200$ $-6.593$
Dropout $-2.400$ $-3.446$ $-2.047$ $+2.590$ $-1.180$ $-0.180$
DropMessage $-3.600$ $-1.125$ $-0.302$ $+0.850$ $-0.460$ $+1.126$
To test the dropping methods on real-world datasets, we use the GCN, GIN [85] and GAT [81] architectures – GCN and GIN satisfy the model assumptions made in all the theoretical results presented in Section 3, while GAT does not satisfy any of them, since the attention scores are computed as a function of all the node representations. Therefore, GCN, GIN and GAT together provide a broad representation of different MPNN architectures. We present the results for GCN and GIN in the main text, since these models were used as baselines in a majority of works on alleviating over-squashing [8, 44, 79, 35, 4, 5, 67]; the results for GAT are reported in Table 6.
For each dataset--model--dropout combination, we perform 20 independent runs to find the best performing dropout configuration; results are reported in Table 9. We then perform a t-test to assess whether dropout improves performance, using 50 samples from the NoDrop model ($q=0$) and 50 samples from the best performing dropout configuration. [^3] The t-test assumes that both samples are drawn from normal distributions – all Shapiro-Wilk tests for non-normality of samples [74] failed at 90% confidence. In this section, we report the p-values of the tests, and in Table 7, we report the effect sizes as Hedges’ $g$ statistic [39].
@@ 5.2 Real-world Datasets
Node-classification. Although determining whether a task requires modelling LRIs can be challenging, understanding the structure of the datasets can provide important insight. For example, homophilic datasets have local consistency in node labels, i.e. nodes closely connected to each other have similar labels. On the other hand, in heterophilic datasets, nearby nodes often have dissimilar labels. Since DropEdge-variants increase the sensitivity of a node’s representations to its immediate neighbors, and reduce its sensitivity to distant nodes, we expect it to improve performance on homophilic datasets but harm performance on heterophilic ones; such a setup was also used in Topping et al. [79]. In this work, we use Cora [57], CiteSeer [31] and PubMed [62] as representatives of homophilic datasets [98, 52], and Squirrel, Chameleon and TwitchDE [69] to represent heterophilic datasets [52]. The networks’ statistics are presented in Table 4, where we can note the remarkably lower homophily measures of heterophilic datasets.
The results are presented in 1(a). It is clear to see that dropout significantly improves test performance on homophilic datasets – with $40/54\approx 74\%$ cases performing better than the corresponding NoDrop baseline – indicating that these methods are indeed beneficial in tackling short-range tasks. On the other hand, with the heterophilic datasets, the improvement is insignificant. Rather, in most ($45/54\approx 83\%$) cases, the best dropout configuration performs worse than the NoDrop baseline. This suggests that the dropping methods harm generalization in long-range tasks by forcing models to overfit to short-range signals (see Section F.3 for supporting evidence).
Graph-classification. Several graph classification datasets have also been identified as long-range tasks, like the molecular networks datasets Mutag [17], Proteins [19] and Enzymes [9], and the social networks datasets Reddit, IMDb and Collab [87]. These datasets have also been used for evaluation in previous works on over-squashing, including Karhadkar et al. [44], Black et al. [8].
The results are shown in 1(b), where we observe that dropout methods generally have insignificant effects on model performance, and often even a non-positive effect ($67/108\approx 62\%$ cases). Notably, the p-values are lower as compared to those recorded for heterophilic datasets in 1(a), i.e. higher evidence for efficacy of dropping methods. We conjecture that over-squashing may have limited impact on model performance in graph-level tasks since the aggregation module eventually mixes information from distant nodes for computing graph-level representations.
5.3 Evaluating DropSens
Figure 3: Relative change in test-time performance of a GCN using DropSens, compared to the baseline DropEdge, on real-world datasets from Section 5.2.
We start by comparing DropSens with DropEdge on real-world datasets from Section 5.2, illustrating how algorithms can be readily adapted for better suitability at modelling LRIs. In Figure 3, we present the relative change in error rate ($1-\text{Acc}$) of DropSens w.r.t. DropEdge, with GCN as the base model. It is clear to observe a uniform improvement in the performance on long-range tasks, suggesting that addressing over-squashing using DropSens can enhance the effectiveness of GCNs.
Table 2: Performance of GCN with graph rewiring methods. First, second, and third best results are coloured.
(a) Node-classification tasks – results for other methods taken from [79, Table 2].
Rewiring Cora CiteSeer PubMed Chameleon Squirrel Actor
None $81.89$ $72.31$ $78.16$ $41.33$ $30.32$ $23.84$
Undirected - - - $42.02$ $35.53$ $21.45$
+FA $81.65$ $70.47$ 79.48 $42.67$ $36.86$ $24.14$
DIGL (PPR) 83.21 73.29 $78.84$ $42.02$ $33.22$ $24.77$
DIGL + Undirected - - - $42.68$ $32.48$ 25.45
SDRF 82.76 72.58 79.10 42.73 37.05 28.42
SDRF + Undirected - - - 44.46 37.67 28.35
DropSens 84.98 73.35 84.30 53.01 41.32 $22.38$
(b) Graph-classification tasks – results for other methods from [8, Table 1].
Rewiring Mutag Proteins Enzymes Reddit IMDb Collab
None $72.15$ $70.98$ $27.67$ $68.26$ 49.77 $33.78$
Last FA $70.05$ 71.02 $26.47$ $68.49$ $48.98$ $33.32$
Every FA $70.45$ $60.04$ $18.33$ $48.49$ $48.17$ 51.80
DIGL 79.70 $70.76$ 35.72 76.04 64.39 54.50
SDRF $71.05$ $70.92$ 28.37 $68.62$ $49.40$ $33.45$
FoSR 80.00 73.42 $25.07$ 70.33 $49.66$ $33.84$
GTR 79.10 72.59 $27.52$ $68.99$ 49.92 $33.05$
DropSens $70.20$ $70.61$ 36.01 77.44 $49.32$ 61.68
We now benchmark DropSens against state-of-the-art graph-rewiring techniques designed specifically to tackle over-squashing (see Section A.3 for their descriptions). We train a GCN on node classification tasks, following the setup in [79], and both GCN and GIN on graph classification tasks, following Karhadkar et al. [44], Black et al. [8]. The results with GCN are reported in 2(a) and 2(b), where we find that DropSens outperforms other methods in node classification tasks, and performs competitively in graph classification tasks. In addition to superior performance, another advantage of DropSens over the other methods is that it significantly reduces the number of messages being propagated, thereby tackling the problem of over-smoothing and increasing training speed.
The results with GIN are presented in Section G.3, where we observe that DropSens does not perform competitively – unsurprising, since DropSens was specifically designed to work with GCN’s message-passing scheme.
@@ 6 Conclusion
There exists an important gap in our understanding of several algorithms designed for training deep GNNs – while their positive effects on model performance have been well-studied, making them popular choices for training deep GNNs, their evaluation has been limited to short-range tasks. This is rooted in a key assumption: that if a deep GNN is trainable, it must also be capable of modelling LRIs. As a result, potential adverse effects of these algorithms on capturing LRIs have been overlooked. Our results challenge this assumption – we theoretically and empirically show that DropEdge- and Dropout-variants exacerbate the over-squashing problem in deep GNNs, and degrade performance on long-range tasks. This highlights the need for a more comprehensive evaluation of common training practices for deep GNNs, with special emphasis on their capacity to capture LRIs. This is crucial for building confidence in their use beyond controlled benchmarks.
Limitations. While our theoretical analysis successfully predicts how DropEdge-variants affect test performance on short-range and long-range tasks, it is based on several simplifying assumptions on the message-passing scheme. These assumptions, although standard in the literature, limit the generalizability of our conclusions to other architectures, including ResGCNs [51], GATs [81], and Graph Transformers [83]. Additionally, an important limitation of DropSens is that it requires an architecture-specific alteration to the edge-dropping strategy, which is not practical in general. As also mentioned in Section 4, we did not intend to introduce DropSens as a benchmark, but rather to demonstrate how methods designed for alleviating over-smoothing can be readily adapted to simultaneously control over-squashing.
Future Directions. Currently, real-world datasets are classified as short- or long-range tasks based on extensive model training [4] or weak proxy measures like node homophily [79]. Developing a reliable measure of information mixing in the ground-truth data could greatly benefit the research community. Such a measure would enable more precise identification of short-, intermediate- and long-range tasks, improving evaluation and benchmarking. Another interesting direction is to investigate the significance of over-squashing in graph-level tasks, where the aggregation module of MPNNs enables some mixing of information from distant nodes. To the best of our knowledge, Giovanni et al. [33] is the only work that directly addresses this question, offering strong theoretical insights. However, empirical validation of these effects remains limited.
Appendix
[…]